Let S = { a + ib / a , b ∈ Z ,b is even } Show that S is a
subring of Z[i] , but not an ideal of Z[i]
1
Expert's answer
2020-09-22T16:18:14-0400
"Soluton"

Step 1

A ring R is a set with two binary operations, addition and multiplication, satisfying several properties: R is an Abelian group under addition, and the multiplication operation satisfies the associative law


"a(bc)=(ab)c"

and distributive laws

"a(b+c)=ab+ac"


and

"(b+c)a=ba+ca"

for every

"a,b,c \in R"

The identity of the addition operation is denoted 0. If the multiplication operation has an identity, it is called a unity. If multiplication is commutative, we say that R is commutative.


A subring of a ring R is a subset


"S \subseteq R"

that is a ring under the operations of R.

Let A be a subring of a ring R. If, for every "a \in A" and "r \in R"


"ar \in A"

and

"ra \in A,"


we say that A is an ideal of R.

The ring "Z[i]"  of Gaussian integers is defined as

"{a+bi/ a,b \in Z}"

under ordinary addition and multiplication of complex numbers.


Step 2

The Subring Test states that a nonempty subset S of a ring R is a subring if and only if it is closed under subtraction and multiplication; that is, if

"a,b \in S,"


then,

"a-b, ab \in S"

Let

"a+bi, c+di \in S"


Step 3

"(a+bi)-(c+di)= (a-c)+(b-d)i"

Since

"a+bi, c+di \in S"

both b and d are even. Hence "b-d"  is even, and

"(a+bi)-(c+di) \in S"


Step 4

Similarly,

"(a+bi)( c+di) =ac+adi+bci-bd \implies (ac-bd)(ad+bc)i"


Step 5

Since b and d are even, "ad" and "bc", are even, and so is "ad+bc". It follows that

"(a+bi)( c+di) \in S"


Step 6

However,

"I \in S"

and

"Ii=i \notin S."

We conclude that S is not an ideal of "Z[i]."